OpenStudy (lgbasallote):

if \(A = \{\emptyset, \{\emptyset\}, \{\emptyset, \{\emptyset\}\}\}\) what is |A|

OpenStudy (zugzwang):

I think it's 3

OpenStudy (lgbasallote):

why so?

OpenStudy (zugzwang):

I counted them?

OpenStudy (lgbasallote):

but they're null....

OpenStudy (zugzwang):

It's set of sets (of sets), it's actually making me dizzy o.O

OpenStudy (zugzwang):

Let U = {} \[V = P(U) = \left\{ \emptyset \right\}\] \[W = P(V) = \left\{ \emptyset, \left\{ \emptyset \right\} \right\}\] \[X = P(W) = \left\{ \emptyset, \left\{ \emptyset \right\}, \ \left\{ \left\{ \emptyset \right\} \right\}, \ \left\{ \emptyset, \ \left\{ \emptyset \right\} \right\} \right\}\] |X| = 4 Now it's just a matter of taking away {{Ø}} making the cardinality 3

OpenStudy (lgbasallote):

....you made it more complicated than it's supposed to be.....

OpenStudy (zugzwang):

Well, my simpler answer was... I counted them :)